# Next-Gen Interconnection Systems with Compute Express Link: a Comprehensive Survey

### v1

@@ III-A CXL Based Memory Expansion
The most common application scenarios is using CXL Type3 Device for Memory Expansion, which breaks the physical constraints of a single machine, allows hosts to access the huge memory on the device through more efficient Load/Store semantics and with lower latency compared with other memory expansion approaches(NVMe, RDMA, etc.). Currently, there is a lot of work on using CXL to alleviate the memory bottleneck faced by memory intensive applications, such as ANN, Large Recommendation Systems, Graph Processing, and In-Memory Databases.
III-A1 Tiered Memory System
Tiered Memory is an OS-level mechanism for managing heterogeneous memory, especially CXL-enabled systems.
@@ III-A1 Tiered Memory System
The current tiered CXL memory systems[7], [8], [40, 41], [42], [43], represented by the TPP[7], efficiently places hot/cold pages in appropriate memory tiers, ensuring performance-critical pages are in fast local memory while offloading less critical ones to slower CXL-Memory. It operates transparently without application-specific knowledge and has been integrated into the Linux kernel.
@@
III-A2 Flash Memory as a Backend
@@
Works represented by the CXL-SSD[14, 46, 47] poses three major challenges of using flash memory: granularity mismatch between CPU and SSD, which leading to traffic and write amplification, significant latency compared to DRAM, and limited endurance due to a finite number of program/erase cycles.
@@
III-A3 Application-Specific Optimizations
@@
Deep learning(DL) workloads also have significant memory requirements, especially Large Recommendation System and Large Language Model. Moiz Arif et al. are the first to use CXL for DL workload‘s memory extension. They propose DeepMemoryDL [49], which manages the allocation of additional CXL-based memory and provides intelligent prefetching and caching mechanisms for DL workloads. Dong Xu et al.[50] improve the existing tensor-offloading approach. They use CXL to build a cache coherence domain between CPU memory and accelerator memory, slightly extend CXL to support an update cache-coherence protocol and avoiding unnecessary data transfers. Miryeong Kwon et al. propose TRAININGCXL [21] that can efficiently process large-scale recommendation datasets in CXL memory pool while making training fault tolerant with low overhead. Haifeng Liu et al. introduce ReCXL[16], which utilizes near-memory processing for scalable, efficient recommendation model training. Yupeng Tang et al. explore leveraging CXL memory to store KV cache in LLM serving [26, 51]. Since DL workloads often rely on heterogeneous computing systems, we will detail related works in section III-C and section V.
@@
III-B CXL Based Disaggregated System
Based on the memory expansion on single-machine, CXL 2.0 proposes a solution for Memory Pooling with the CXL Switch. In data centers, servers will be decomposed into memory modules (such as DRAM) and compute modules (such as CPUs), and connected with CXL Switch. CXL Switch organizes all the memory modules into a memory pool and exposes it to the compute modules. The compute modules can dynamically request and release memory modules based on their demands, effectively improving the overall memory utilization efficiency in the data center. In the future, the data center will become a fabric composed of a large number of modular components, leads to Composable Infrastructure.
III-B1 Architecture of Pooling System
In 2022, researchers from KAIST proposes DirectCXL[12], a software-hardware co-designed CXL memory pooling system. DirectCXL is primarily built upon CXL 2.0 and can be considered as one of the earliest publicly disclosed CXL memory pooling prototypes based on FPGA. In DirectCXL, a single CXL Switch connects multiple CXL memory modules and multiple processors to form a memory pool and computation pool. They reveal that CXL‘s‘ latency mainly comes from PCIe bus, not related to the transmission size but CPU cache.
They expanded DirectCXL by incorporating the features of CXL3.0 [60], upgrading it from a memory pool system to a more fine-grained and larger-scale composable system. They integrated the Multi-Header Logical Device (MH-LD) and Dynamic Capacity Device (DCD) features into DirectCXL, providing more fine-grained memory management and sharing capabilities. Additionally, they used a port-number-based routing mechanism to extend the CXL topology, allowing the integration of more CXL switches and devices, forming a large-scale, multi-level switch structure that can theoretically support up to 4096 devices and/or hosts.
Inspired by CXL, Mohammad Ewais et al. discusses the CXL based Disaggregated Datacenter (DDC)[13]. They believe that CXL will help make the vision of a disaggregated datacenter a reality. They propose a system model with the following key elements: Compute Units, Caching, Memory Nodes, Virtual Memory, Interconnects and Networking, Storage, Management. They also discusses the limitations of existing research, such as the lack of a reasonable explanation for architectural choices, the convenience of design and adoption, latency issues, and the complexity of data sharing.
Unlike the physical memory pooling represented by DirectCXL, Emmanuel Amaro et al. from VMware propose a new memory disaggregation architecture based on Logical Memory Pools (LMPs) [61]. Compared to physical memory pooling, LMPs create memory pools by partitioning a portion of the local memory on each server, rather than using a dedicated memory device. The advantage of this approach is lower cost, support for near-memory computing without additional hardware, and more flexibility in design. The evaluation show that LMPs can execute workloads that are not feasible in physical memory pools, and their performance is also better due to faster access speeds.
Ming Liu proposes a new computing paradigm called Fabric-Centric Computing (FCC)[62]. This approach addresses the challenges of emerging memory fabrics and composable infrastructures by treating the memory fabric as a first-class citizen for instance, orchestration, and reclamation of computing resources. Liu discusses the differences between memory fabrics and communication fabrics, including the synchronous execution model, diverse memory node types, dependence on PCIe routing performance, fast execution engine context switching, and the introduction of passive fault domains. Based on these characteristics, Liu proposes the design principles of FCC, including data movement as a managed service, host-assisted memory node type-aware data structures, idempotent tasks and hardware-assisted scalable offloading, as well as a fabric-centric arbitrator implemented over dedicated channels.
III-B2 Interconnection in Pooling System
In disaggregated systems, the distance between storage and computing resources (and the resulting latency) often becomes a bottleneck for system performance.
Andreas Geyer et al. also evaluated the impact of interconnect on disaggregated systems[64]. The paper focuses on discussing RDMA and CXL, which allow dynamic sharing of hardware resources between different servers. they systematically evaluated the impact of local Ultrapath Interconnect (UPI), RDMA over InfiniBand, and memory attached via CXL over PCIe on disaggregated systems. The results show that the performance of CXL is comparable to UPI and better than asynchronous RDMA.
The researchers from UCAS, proposes CXL over Ethernet[66]. By using RDMA to transmit CXL access requests, it overcomes the shortcomings of native CXL-based disaggregated systems, which are limited to the rack scale. The CXL memory requests can be directly sent to the FPGA, which then interacts with the remote memory through its network interface. This work demonstrates that it is possible to combine the advantages of RDMA and native CXL technologies.
The collaborative work between the Ming Liu and Zeke Wang from Zhejiang University propose rPCIeBench [68], a hardware-software co-designed benchmark framework for systematically evaluating the performance of Routable PCIe (CXL), which is the key interconnect for building new composable infrastructures. They first analyzes the communication characteristics of the Routable PCIe and compares it with local PCIe. It then uses rPCIeBench to delve into the traffic coordination behavior within the fabric, and proposes three interesting findings: Approximate Max-Min Bandwidth Partition, Fast End-to-End Bandwidth Synchronization, and Interference-free Orthogonal Paths. Finally, they translate these insights into traffic coordination rules and develop an edge-constraint relaxation algorithm to estimate the transfer performance of PCIe flows on the shared fabric.
III-B3 Cost of Pooling System
Researchers from Google[69] have raised concerns about the potential application of CXL memory pooling in data centers and cloud systems. The paper points out three main issues that hinder the practicality of CXL memory pooling: cost, complexity, and utility. First, the cost of CXL memory pooling may outweigh the savings from reduced RAM. Second, the actual latency of CXL is higher than main memory, which requires extensive rewriting of networked applications, increasing software complexity. Finally, by analyzing production traces from Google and Azure Cloud, it was found that modern servers are large enough for most virtual machines, and simple virtual machine packing algorithms hardly leave any unused memory, which undermines the main motivation for memory pooling.
Contrary to the opinion of Google, Researchers from Microsoft argues that CXL-based memory pooling has significant cost advantages, and discusses some key design considerations for implementing a CXL-based disaggregated memory system in cloud computing scenarios[44]. They analyzes the critical design constraints for cloud service providers deploying CXL memory pools (memory pool access performance, compatibility with virtualization, memory pool system topology, etc.) as well as observations from actual deployments, and points out configuration examples with significant return on investment. They also compares the costs of memory pools with different pool sizes and CXL device types.
III-C Near-Memory Processing
@@
Near-Memory Processing(NMP) is a promising research topic in the field of computer architecture. CXL Memory Device does not have computing capability, the data must go through an costly data movement to be processed. Pioneer researchers are working to introduce near-memory computing in CXL systems, offloading some computing functions to the CXL memory devices, so that the computation can be executed on the memory side, avoiding data movement.
III-C1 Workload-Customized Near-Memory Process
@@
III-C2 Generalized Near-Memory Process
Joonseop Sim et al. from SK Hynix proposed a solution called CMS[73] to address the bandwidth limitation of CXL devices. CMS integrates NDP cores in the CXL device, which reduces the data movement through the CXL interface. CMS also proposes methods to improve CMS bandwidth utilization through hardware logic load balancers and MAC block optimizations.
@@ III-C2 Generalized Near-Memory Process
Hyungkyu Ham et al. from POSTECH proposed M2NDP (Memory-Mapped NDP) [75], designed specifically for CXL memory expanders. M2NDP achieves low-cost, general-purpose NDP functionality by combining memory-mapped functions (M2func) and memory-mapped micro-threading (M2thr). M2func utilizes the unmodified CXL.mem protocol to enable lightweight communication between the host and CXL devices, avoiding the high overhead of traditional PCIe/CXL.io solutions. M2thr introduces micro-threads, a lightweight thread with minimal register allocation, allowing a large number of micro-threads to execute concurrently on low-cost NDP units.
Jon Hermes et al. proposed UDON[76], exploring the potential and benefits of offloading tasks to general-purpose cores on CXL memory devices. They simulates a CXL type-2 device (a device with general-purpose compute cores) on an Arm AArch64 dual-socket NUMA system. The study finds that by intelligently partitioning the machine learning inference model for computation offloading, up to 90% of the data can be placed in CXL memory with only a 20% performance loss.
IV Distributed Shared Memory
Unlike memory pooling proposed in CXL2.0, CXL3.0 introduces Memory Sharing. A memory module is occupied by a single compute module in CXL2.0, whereas memory sharing allows a memory module to be shared by multiple compute modules (potentially across multiple servers). CXL3.0 provides a hardware cache coherence mechanism, effectively avoids the major performance bottleneck of traditional distributed shared memory systems: the network-based cache coherence mechanism, and opens up new opportunities.
IV-A Distributed Shared Memory with CXL 3.0
IV-B Communication via Shared CXL Memory
In CXL-SHM[18], Zhang et al. also proposes a pass-by-reference RPC framework that improves the efficiency of communication by passing object references (rather than object values) in shared CXL memory.
On the basis of CXL-SHM[18], researchers from Alibaba proposed HydraRPC[78], an RPC system based on shared CXL memory. HydraRPC introduces a novel non-cached sharing mechanism. In data reading, it sets specific memory areas as uncachable to ensure the data is transferred directly between CXL HDM and the processor, bypassing the CPU cache. In data writing, HydraRPC uses specific CPU instructions to ensure that data is written directly to the CXL HDM, rather than just to the local cache. This synchronizes non-temporal accesses using memory load/store fences, ensuring data consistency.
IV-C Failure Tolerance
In CXL-SHM[18], Zhang et al. uses a special era-based non-blocking algorithm to maintain reference counting, effectively solving the synchronization, memory leaks, double release, and dangling pointer issues in distributed memory management. Even if some clients fail unexpectedly due to process crashes or machine failures, the system can automatically recover, ensuring the stability and reliability.
Steven Swanson et al. from UCSD, firstly propose a failure model that classifies and handles Data Failures and Process Failures in CXL shared memory [80]. First, the researchers discuss data failures, which occur when the CXL memory device becomes unavailable. Next, they explores process failures and compares them to the failure model and programming solutions in persistent memory (PMEM). The researchers note that while PMEM shares some features with CXL, it does not fully address the volatility and low access latency challenges of CXL. This work also propose customized solutions to the specific challenges of CXL systems.
@@ V Unified Memory
The hardware-based cache coherence provided by CXL not only accelerates distributed shared memory, but also enables the Unified Memory across heterogeneous accelerator. In today’s data centers, heterogeneous accelerator (such as GPUs, TPUs, etc.) are gradually becoming the main computing power. However, due to the lack of cache coherence between devices and processors, the main memory and device memory cannot be accessed uniformly, and data interaction between devices requires the CPU to act as a data transfer relay, which has become a performance bottleneck in heterogeneous computing systems. The hardware-based cache coherence allows heterogeneous devices to uniformly access data on each other’s memories, which not only avoid unnecessary data copying, improve the computing efficiency of the CPU, but also extend memory capacity for heterogeneous devices.
@@ V-B Unified Memory with Direct CXL Access
In 2022, Anthony M Cabrera et al. designed and analyzed the performance model of CXL devices in heterogeneous computing systems with existing hardware [82]. Since CXL device prototypes were not yet available, they developed a benchmark application called DecEval to simulate the collaborative work of GPUs and FPGAs using existing software stacks and frameworks. The CXL performance model was based on high-level expectations of how CXL would operate, including a theoretical analysis that compares the performance of CXL and PCIe. They used existing PCIe models as a comparison baseline and evaluated the CXL model based on measurement data collected from the benchmark application.
@@
The team leads by Myoungsoog Jung, delves deeper into how to use CXL to extend and share memory for heterogeneous accelerators in a unified memory manner[20]. They integrats CXL into the GPU. The GPU contains multiple CXL root ports, each equipped with a CXL controller that support DRAM or SSD. Through speculative reads and deterministic write mechanisms, allowing the GPU to predict and prefetch data, while also handling the latency of write operations. The GPU’s system bus and memory mapping allow the GPU to directly issue memory requests, which are handled by the CXL root port, without the need for CPU involvement.
@@
Moiz Arif et al. discussed the resource contention issues that arise when accelerating GPU-based workloads through CXL[84]. They believe that the contention mainly occurs in two aspects. First, the PCIe bandwidth: when multiple GPUs access CXL memory simultaneously, it can saturate the PCIe bandwidth. Second, the memory utilization: traditional memory allocation and task scheduling can lead to imbalanced memory utilization among multiple GPUs They propose a memory allocation strategy that considers the memory requirements of each job and the available memory resources. Based on the current memory usage of the jobs and the overall memory demand of the system, the strategy can also dynamically adjust the memory allocation to reduce memory contention. The strategy also considers the bandwidth limitations and access latency of the PCIe and CXL to avoid bandwidth contention.
@@
VI Future Research

### v2

@@ III Memory Expansion
As shown in figure 2, the most common application scenarios is using CXL Type3 Device for Memory Expansion, which breaks the physical constraints of a single machine, allows hosts to access the huge memory on the device through Load/Store semantics and with lower latency compared with other memory expansion approaches(NVMe, RDMA, etc.). Currently, there is a lot of work on using CXL to alleviate the memory bottleneck faced by memory intensive applications, such as ANN, Large Recommendation Systems, Graph Processing, and In-Memory Databases.
Figure 2: CXL-based Memory Expansion. It is based on tiered memory, which places data in appropriate memory tiers. In order to avoid the overhead of data movement between tiers, researches introduce near-memory processing.
III-A Tiered Memory System
@@
Although CXL Memory has a larger capacity, however, its latency is much higher than DRAM, which causes performance degradation. Tiered Memory is an OS-level mechanism for leveraging heterogeneous memory, especially CXL-enabled systems. The current tiered CXL Memory systems[9], [10], [35, 36], [37], [38], represented by the TPP[9], efficiently places hot/cold pages in appropriate memory tiers, ensuring performance-critical pages are in fast DRAM memory while offloading less critical ones to slower CXL Memory. It operates transparently without application-specific knowledge and has been integrated into the Linux kernel.
@@
III-B Flash Memory as a Backend
@@
Works represented by the CXL-SSD[39, 40],[41] poses three major challenges of using flash memory: granularity mismatch between CPU and SSD, which leading to traffic and write amplification, significant latency compared to DRAM, and limited endurance due to a finite number of program/erase cycles.
@@
III-C Application-Specific Optimizations
@@
Deep learning(DL) workloads also have significant memory requirements, especially Large Recommendation System and Large Language Model. Moiz Arif et al. are the first to use CXL for DL workload‘s memory extension. They propose DeepMemoryDL [42], which manages the allocation of additional CXL-based memory and provides intelligent prefetching and caching mechanisms for DL workloads. Dong Xu et al.[80] improve the existing tensor-offloading approach. They use CXL to build a unified memory between main memory and accelerator memory, enabling accelerators to use main memory directly, addressing the storage challenge of deep learning workloads Miryeong Kwon et al. propose TRAININGCXL [16] that can efficiently process large-scale recommendation datasets in CXL memory pool while making training fault tolerant with low overhead. Haifeng Liu et al. introduce ReCXL[14], which utilizes near-memory processing for scalable, efficient recommendation model training. Yupeng Tang et al. explore leveraging CXL memory to store KV cache in LLM serving [25, 43]. Since DL workloads often rely on heterogeneous computing systems, we will detail related works in section IV and section V.
@@
IV Near-Memory Processing
@@
The reason for the high latency of accessing CXL memory is that CXL Type3 Device itself does not have computing capability, the data must go through a costly movement from device to processor to be processed. A possible solution is Near-Memory Processing(NMP). Pioneer researchers are working to introduce NMP to CXL systems, offloading some computing functions to the CXL memory devices, so that the computation can be executed on the memory side (as shown in figure 2), avoiding data movement.
IV-A Workload-Customized Near-Memory Process
@@
IV-B Generalized Near-Memory Process
Joonseop Sim et al. from SK Hynix propose a solution called CMS[47] to address the bandwidth limitation of CXL devices. CMS integrates NDP cores in the CXL device, which reduces the data movement through the CXL interface. CMS also proposes methods to improve CMS bandwidth utilization through hardware logic load balancers and MAC block optimizations.
@@ IV-B Generalized Near-Memory Process
Hyungkyu Ham et al. from POSTECH propose M2NDP (Memory-Mapped NDP) [49], designed specifically for CXL memory expanders. M2NDP achieves low-cost, general-purpose NDP functionality by combining memory-mapped functions (M2func) and memory-mapped micro-threading (M2thr). M2func utilizes the unmodified CXL.mem protocol to enable lightweight communication between the host and CXL devices, avoiding the high overhead of traditional PCIe/CXL.io solutions. M2thr introduces micro-threads, a lightweight thread with minimal register allocation, allowing a large number of micro-threads to execute concurrently on low-cost NDP units.
Jon Hermes et al. propose UDON[50], exploring the potential and benefits of offloading tasks to general-purpose cores on CXL memory devices. They simulate a CXL type-2 device (a device with general-purpose compute cores) on an Arm AArch64 dual-socket NUMA system. This study finds that by intelligently partitioning the machine learning inference model for computation offloading, up to 90% of the data can be placed in CXL memory with only a 20% performance loss.
@@ V Unified Memory
Figure 3: Unified Memory. The hardware cache coherence provided by CXL organises the device-attached memory and host-attached memory into a unified memory space.
CXL system include not only processors and memory devices, but also heterogeneous accelerators (Type2 Device, such as GPUs, TPUs, etc.). Unlike memory devices, accelerators have both compute units and memory. However, due to the lack of cache coherence between accelerators and processors, the main memory and device memory cannot be accessed uniformly. The data interaction between devices requires the CPU to act as a data transfer relay, which has become a performance bottleneck in heterogeneous computing systems.
As shown in figure 3, CXL provides hardware-based cache coherence to enable the Unified Memory across accelerators and processors. It allows accelerators to uniformly access data on each other’s memories, which not only avoid unnecessary data copying, improve the computing efficiency of the CPU, but also extend memory capacity for heterogeneous devices.
@@ V-B Unified Memory with Direct CXL Access
In 2022, Anthony M Cabrera et al. design and analyze the performance model of CXL devices in heterogeneous computing systems with existing hardware [52]. Since CXL device prototypes are not yet available, they develop a benchmark application called DecEval to simulate the collaborative work of GPUs and FPGAs using existing software stacks and frameworks. The CXL performance model is based on high-level expectations of how CXL would operate, including a theoretical analysis that compares the performance of CXL and PCIe. They use existing PCIe models as a comparison baseline and evaluate the CXL model based on measurement data collected from the benchmark application.
@@
The team led by Myoungsoog Jung delves deeper into how to use CXL to extend and share memory for heterogeneous accelerators in a unified memory manner[15]. They integrate CXL into the GPU. The GPU contains multiple CXL root ports, each equipped with a CXL controller that support DRAM or SSD. Through speculative reads and deterministic write mechanisms, allowing the GPU to predict and prefetch data, while also handling the latency of write operations. The GPU’s system bus and memory mapping allow the GPU to directly issue memory requests, which are handled by the CXL root port, without the need for CPU involvement.
@@
Moiz Arif et al. discuss the resource contention issues that arise when accelerating GPU-based workloads through CXL[54]. They believe that the contention mainly occurs in two aspects. First, the PCIe bandwidth: when multiple GPUs access CXL memory simultaneously, it can saturate the PCIe bandwidth. Second, the memory utilization: traditional memory allocation and task scheduling can lead to imbalanced memory utilization among multiple GPUs They propose a memory allocation strategy that considers the memory requirements of each job and the available memory resources. Based on the current memory usage of the jobs and the overall memory demand of the system, the strategy can also dynamically adjust the memory allocation to reduce memory contention. The strategy also considers the bandwidth limitations and access latency of the PCIe and CXL to avoid bandwidth contention.
Dong Xu et al.[80] improve the existing tensor-offloading approach. They use CXL to build a cache coherence domain between CPU memory and accelerator memory, slightly extend CXL to support an update cache-coherence protocol and avoiding unnecessary data transfers.
@@
VI CXL Based Disaggregated System
Figure 4: CXL Based Disaggregated System[2]. It breaks down the physical isolation between servers and consolidates the data center into a hyper-node through resource pooling.
Based on the memory expansion on single-machine, CXL 2.0 proposes a solution for Memory Pooling with the CXL Switch, as shown in figure 4.
In data centers, servers will be decomposed into memory modules (such as DRAM) and compute modules (such as CPUs), and connected with CXL Switch. CXL Switch organizes all the memory modules into a memory pool and exposes it to the compute modules. The compute modules can dynamically request and release memory modules based on their demands, effectively improving the overall memory utilization efficiency in the data center. In the future, the data center will become a fabric composed of a large number of modular components, leads to Composable Infrastructure.
VI-A Architecture of Pooling System
In 2022, researchers from KAIST propose DirectCXL[19], a software-hardware co-designed CXL memory pooling system. DirectCXL is primarily built upon CXL 2.0 and can be considered as one of the earliest publicly disclosed CXL memory pooling prototypes based on FPGA. In DirectCXL, a single CXL Switch connects multiple CXL memory modules and multiple processors to form a memory pool and computation pool. They reveal that CXL‘s‘ latency mainly comes from PCIe bus, not related to the transmission size but CPU cache.
They expand DirectCXL by incorporating the features of CXL3.0 [56], upgrading it from a memory pool system to a more fine-grained and larger-scale composable system. They integrate the Multi-Header Logical Device (MH-LD) and Dynamic Capacity Device (DCD) features into DirectCXL, providing more fine-grained memory management and sharing capabilities. Additionally, they use a port-number-based routing mechanism to extend the CXL topology, allowing the integration of more CXL switches and devices, forming a large-scale, multi-level switch structure that can theoretically support up to 4096 devices and/or hosts.
Inspired by CXL, Mohammad Ewais et al. discuss the CXL based Disaggregated Datacenter (DDC)[20]. They believe that CXL will help make the vision of a disaggregated datacenter a reality. They propose a system model with the following key elements: Compute Units, Caching, Memory Nodes, Virtual Memory, Interconnects and Networking, Storage, Management. They also discuss the limitations of existing research, such as the lack of a reasonable explanation for architectural choices, the convenience of design and adoption, latency issues, and the complexity of data sharing.
Unlike the physical memory pooling represented by DirectCXL, Emmanuel Amaro et al. from VMware propose a new memory disaggregation architecture based on Logical Memory Pools (LMPs) [57]. Compared to physical memory pooling, LMPs create memory pools by partitioning a portion of the local memory on each server, rather than using a dedicated memory device. The advantage of this approach is lower cost, support for near-memory computing without additional hardware, and more flexibility in design. The evaluations show that LMPs can execute workloads that are not feasible in physical memory pools, and their performance is also better due to faster access speeds.
Ming Liu proposes a new computing paradigm called Fabric-Centric Computing (FCC)[58]. This approach addresses the challenges of emerging memory fabrics and composable infrastructures, treating the memory fabric as a first-class citizen for instance, orchestration, and reclamation of computing resources. Liu discusses the differences between memory fabrics and communication fabrics, including the synchronous execution model, diverse memory node types, dependence on PCIe routing performance, fast execution engine context switching, and the introduction of passive fault domains. Based on these characteristics, Liu proposes the design principles of FCC, including data movement as a managed service, host-assisted memory node type-aware data structures, idempotent tasks and hardware-assisted scalable offloading, as well as a fabric-centric arbitrator implemented over dedicated channels.
VI-B Interconnection in Pooling System
In disaggregated systems, the distance between storage and computing resources (known as the Access Latency) often becomes a bottleneck to overall performance.
Andreas Geyer et al. also evaluate the impact of interconnect on disaggregated systems[60]. The paper focuses on discussing RDMA and CXL, which allow dynamic sharing of hardware resources between different servers. they systematically evaluated the impact of local Ultrapath Interconnect (UPI), RDMA over InfiniBand, and memory attached via CXL over PCIe on disaggregated systems. The results show that the performance of CXL is comparable to UPI and better than asynchronous RDMA.
The researchers from UCAS, propose CXL over Ethernet [62]. By using RDMA to transmit CXL access requests, it overcomes the shortcomings of native CXL-based disaggregated systems, which are limited to the rack scale. The CXL memory requests can be directly sent to the FPGA, which then interacts with the remote memory through its network interface. This work demonstrates that it is possible to combine the advantages of RDMA and native CXL technologies.
The collaborative work between Ming Liu and Zeke Wang from Zhejiang University propose rPCIeBench [64], a hardware-software co-designed benchmark framework for systematically evaluating the performance of Routable PCIe (CXL), which is the key interconnect for building new composable infrastructures. They first analyzes the communication characteristics of the Routable PCIe and compares it with local PCIe. Then, they use rPCIeBench to delve into the traffic coordination behavior within the fabric, and propose three interesting findings: Approximate Max-Min Bandwidth Partition, Fast End-to-End Bandwidth Synchronization, and Interference-free Orthogonal Paths. Finally, they translate these insights into traffic coordination rules and develop an edge-constraint relaxation algorithm to estimate the transfer performance of PCIe flows on the shared fabric.
VI-C Cost of Pooling System
Researchers from Google[65] have raised concerns about the potential application of CXL memory pooling in data centers and cloud systems. The paper points out three main issues that hinder the practicality of CXL memory pooling: cost, complexity, and utility. First, the cost of CXL memory pooling may outweigh the savings from reduced RAM. Second, the actual latency of CXL is higher than main memory, which requires extensive rewriting of networked applications, increasing software complexity. Finally, by analyzing production traces from Google and Azure Cloud, it is found that modern servers are large enough for most virtual machines, and simple virtual machine packing algorithms hardly leave any unused memory, which undermines the main motivation for memory pooling.
Contrary to the opinion of Google, Researchers from Microsoft argue that CXL-based memory pooling has significant cost advantages, and discuss some key design considerations for implementing a CXL-based disaggregated memory system in cloud computing scenarios[66]. They analyze the critical design constraints for cloud service providers deploying CXL memory pools (memory pool access performance, compatibility with virtualization, memory pool system topology, etc.) as well as observations from actual deployments, and point out configuration examples with significant return on investment. They also compare the costs of memory pools with different pool sizes and CXL device types.
VII Distributed Shared Memory
Figure 5: Communication via Shared CXL Memory
Unlike memory pooling proposed in CXL2.0, CXL3.0 introduces Memory Sharing. A memory module is occupied by a single compute module in CXL2.0, whereas memory sharing allows a memory module to be shared by multiple compute modules, as shown in figure 4. CXL3.0 provides a hardware cache coherence mechanism, effectively avoids the major performance bottleneck of traditional distributed shared memory systems: the network-based cache coherence mechanism, and opens up new opportunities.
VII-A Distributed Shared Memory with CXL 3.0
VII-B Communication via Shared CXL Memory
As shown in figure 5, the shared CXL memory enables efficient synchronisation and communication between distributed processes running on different compute modules. The client process writes a request in the shared memory, and the server process polls the shared memory, detects the request, processes it, and sends response.
In CXL-SHM[22], Zhang et al. also propose a pass-by-reference RPC framework that improves communication efficiency by passing object references (rather than object values) in shared CXL memory.
On the basis of CXL-SHM[22], researchers from Alibaba propose HydraRPC[69], an RPC system based on shared CXL memory. HydraRPC introduces a novel non-cached sharing mechanism. In data reading, it sets specific memory areas as uncachable to ensure that the data is transferred directly between CXL HDM and the processor, bypassing the CPU cache. In data writing, HydraRPC uses specific CPU instructions to ensure that data is written directly to the CXL HDM, rather than just to the local cache. This synchronizes non-temporal accesses using load/store fences, ensuring consistency.
VII-C Failure Tolerance
In CXL-SHM[22], Zhang et al. use a special era-based non-blocking algorithm to maintain reference counting, effectively solving the synchronization, memory leaks, double release, and dangling pointer issues in distributed memory management. Even if some clients fail unexpectedly due to process crashes or machine failures, the system can automatically recover, ensuring stability and reliability.
Steven Swanson et al. from UCSD, firstly propose a failure model that classifies and handles Data Failures and Process Failures in CXL shared memory [71]. First, the researchers discuss data failures, which occur when the CXL memory device becomes unavailable. Next, they explore process failures and compare them to the failure model and programming solutions in persistent memory (PMEM). The researchers note that while PMEM shares some features with CXL, it does not fully address the volatility and low access latency challenges of CXL. This work also proposes customized solutions to the specific challenges of CXL systems.
VIII Future Research